Model-Reference Control

Stanislaw H. Żak ^{∗} May 17, 2000

Abstract

Different schemes involving model-reference control are analyzed.

∗School of Electrical and Computer Engineering, Purdue University, West Lafayette, IN 47907

Figure 1: Adaptive inverse control.

Figure 2: Canceling plant noise and disturbance. 2